OpenStudy (anonymous):

Solve the following system of equations alebraically: 2x+3y=17 3x-2y=-0.5 x= y=

OpenStudy (anonymous):

ok well first u need to find one variable so u do 2x+3y=17 2x=(-3y+17)/2 x=-3/2 y +17/2 3x-2y=-.5 now u substitue so 3(-3/2 y + 17/2) -2y=-.5 -4.5+25.5-2y=-.5 21-2y=-.5 -2y=20.5 y=10.25

OpenStudy (anonymous):

now you got y, you can find x so 2x+3y=17 2x+3(10.25)=17 2x+30.75)=17 2x=-13.25 x=-6.625